# Canonical formalism in the Schwarzschild spacetime

### v2

@@ Abstract
This work constitutes the second part of a series of studies that aim to utilise tools from Hamiltonian mechanics to investigate the motion of an extended body in general relativity. The first part of this work [Refs. [1, 2]] constructed a ten-dimensional, covariant Hamiltonian framework encompassing all the linear-in-spin corrections to the geodesic motion in arbitrary spacetime. This framework was proven to be integrable in the Schwarzschild and Kerr spacetimes, specifically. The present work translates this abstract integrability result into tangible applications for linear-in-spin Hamiltonian dynamics of a compact object in a Schwarzschild spacetime. In particular, a canonical system of coordinates is constructed explicitly, which exploits the spherical symmetry of the Schwarzschild spacetime. These coordinates are based on a relativistic generalization of the classical Andoyer variables of Newtonian rigid body motion. This canonical setup allows us to derive ready-to-use formulae for action-angle coordinates and gauge-invariant Hamiltonian frequencies, which automatically include all linear-in-spin effects. No external parameters or ad hoc choices are necessary, and the framework can be used to find complete solutions by quadrature of generic (bound or unbound), linear-in-spin orbits, including orbital inclination, precession and eccentricity, as well as spin precession. The efficacy of the formalism is demonstrated here in the context of circular orbits with arbitrary spin and orbital precession, with the results validated against known results in the literature.
@@
The main motivation of our efforts comes from extreme mass-ratio inspirals (EMRIs) – the gradual inspiral of a stellar-mass compact object (typically a neutron star or black hole or even an “exotic” compact object) into a supermassive black hole [3]. These binary inspirals are promising sources of gravitational waves for an interferometric detector in space (e.g., the Laser Interferometer Space Antenna [4, 5], as well as DECIGO [6], TianQin, and Taiji [7, 8, 9]) and they will eventually provide an exquisite probe of the strongly gravitating region close to a black hole; for example, the source’s astrophysical parameters, such as the masses and the spin of the larger central black hole, can be measured with unparalleled precision [10, 11]. However, the gravitational-wave signals from EMRIs will be so intricate and sufficiently weak that they can only be extracted from the noise-dominated detector data with modelled waveforms, that match the phase of the signal over its full duration (scaling like the inverse of the mass ratio) across all possible orbital and spin configurations around a central black hole [12]. As a result, the scientific promise described above will be attained only with a highly accurate and efficient model of relativistic two-body systems, built on the detailed understanding of their orbital dynamics in the (multidimansional) EMRI parameter space.
@@
In what follows, we give some more details about the relativistic two-body problem in the extreme mass-ratio regime, and motivate our work by emphasizing the need for a Hamiltonian treatment of the secondary spin.
@@
Over the years, exact numerical relativity simulations [13], analytical approximation schemes [14, 15, 16, 17] and effective models have been devised to tackle the relativistic two-body problem [18, 19, 20, 21, 22, 23], each specializing to some type of source, but also overlapping with the other in their respective domain of validity [24, 25, 26]. Virtually all of them (perhaps a sole exception is the numerical relativity) benefit from gravitation’s unique feature of universality, i.e., the tendency of a body to move on pre-defined paths in the curved geometry of spacetime, regardless of its internal composition. This feature, however, is only an approximation, as is revealed by multipolar expansion schemes [27, 28, 29, 30]. This modelling approach consists in representing a real, extended compact object as a point particle endowed with a finite number of multipoles, encompassing mass, linear and angular momentum, deformability coefficients, among others. The granularity of this description increases with the number of multipoles included, providing a more realistic description of the object.
@@
In this multipolar expansion approach to the two-body dynamics, the associated small parameter is the typical length scale of the object over the typical length scale of the background spacetime, which is therefore tailor-made for modeling asymmetric binaries, in particular EMRIs. At the “zeroth” order, the motion of a smaller compact object is well approximated by a test body moving along a geodesics of the fixed background spacetime. However, this is only a crude approximation. Additional effects arise and lead to corrections of this geodesic motion [29]. These include (i) conservative and dissipative self-field effects of the small object, owing to the non-linearity of general relativity (GR) and the emission of gravitational waves, and (ii) extended-body effects, due to the compact object not being exactly a point, but rather an extended object which can spin and deform. All these effects are important for measuring EMRIs by space-based gravitational-wave detectors, as all of them have an impact of quadratic order in the mass ratio in the waveform phase [17, 31], which is considered to be the minimal accuracy required for parameter estimation [12, 32].
@@
In the present paper (and its foundation [1, 2]), we leave the self-field effects aside (i.e. we neglect the back-reaction of the small object’s gravitational field), and solely focus on the leading order extended-body effect from the linear-in-spin corrections to the geodesic motion. At this linear-in-spin order, although it is specific to GR, the equations that describe the evolution of linear and angular momenta along the particle’s worldline are purely kinematical (no forces or torques) and universal (all bodies move the same way) [30]. Body-specific forces and torques only arise at the next quadrupolar order, with quadratic-in-spin effects [33].
@@
Analytically solving the relevant linear-in-spin evolution equations, now known as the Mathisson-Papapetrou-Tulczyjew-Dixon (MPTD) equations, on an astrophysically relevant Kerr black-hole background has been the subject of many past studies using a whole spectrum of methods. Among them, state-of-the-art methods use the powerful formalism of the Marck tetrad [34, 35] for the spin to solve the MPTD equations directly in a “perturbed geodesic” philosophy. This method is very natural given the quasi-geodesic nature of linear-in-spin dynamics, especially leading to purely analytical results for generic orbits in the Schwarzschild spacetime [36] and simple orbital configurations in Kerr [37, 38, 39, 31, 40]. A hybrid numerical/analytical result exists for more generic orbits [41, 42, 43, 44]. As such, there is a sense in which this Marck formalism offers a simpler implementation than other means of obtaining the solutions to the linear-in-spin equations of motion. Having the solution trajectory, however, is not the only way to understand the linear-in-spin dynamics of the MPTD system. Many other useful characterizations such as integrability of the system, in which the equations of motion can be solved by quadrature for any orbital configurations, may require another approach to the MPTD problem.
@@
A Hamiltonian formulation just offers such an approach. In the test-mass approximation (i.e. with no spin), the powerful machinery of Hamiltonian mechanics has been already proven its value in the context of motion in curved spacetime. A non-exhausted list of examples includes (but not limited to) the Hamilton-Jacobi equation [45], action-angle variables [46, 47], the phase space analysis [48], Hamiltonian frequencies and resonances [46, 49], the first law of mechanics [50, 51, 52], the perturbation (KAM-type) theory [53, 54], Birkhoff normal forms (through Lie series) [55] symplectic integrators [56, 57]. In mathematical point of view, these tools are indeed brought by the unique integrability of Kerr geodesic dynamics as a Hamiltonian system [58].
Nonetheless, we are nowhere near having a Hamiltonian understanding of these Kerr linear-in-spin dynamics as thorough as that of the Kerr geodesics dynamics. Even though Hamiltonian schemes including linear-in-spin effects exist, for example those constructed in Refs. [59, 60] (see also Refs. [1] for detailed references), the existing Hamiltonian frameworks have provided only incomplete proof of the integrability in Kerr spacetime, or have not benefited from the effectiveness of classical tools from Hamiltonian mechanics.
The present paper, and its foundation in Refs. [2, 1], aim at filling this gap.
.2 Summary of main results
@@
The classical multipolar expansion schemes (reviewed in Sec. I.1) assert that this object can be modelled as a point particle endowed with a linear momentum 1-form $p_{a}$ and spin tensor $S^{ab}$. Their evolution along the particle’s worldline $\mathscr{L}$, at linear order in $S^{ab}$ and under the Tulczyjew-Dixon (TD) spin supplementary condition (which fixes a unique representative worldline for the extended body), is given by the following set of MPTD equations
@@
As previously outlined by one of us (PR) in Ref. [1] (hereafter referred to as “Paper I”) the algebraic-differential system in Eq. (1) can be cast into a covariant Hamiltonian formulation on a $10$ dimensional ($10$D) phase space $\mathcal{P}$. Most importantly, Paper I demonstrates that, when applied to a background Kerr spacetime, the resulting Hamiltonian system is integrable in the usual, Arnold-Liouville sense [61, 62].
@@
In the present paper we take on that endeavor, and construct the canonical coordinates of $\mathcal{P}$ explicitly (see also Ref. [63] for other attempts in this direction). Based on these coordinates, we also derive a set of practically useful coordinates: (so-named) action-angle coordinates, which is unique to integrable systems, as demonstrated by the Arnold-Liouville theorem in Ref. [62]. Action-angle coordinates permit the calculation of frequencies associated with oscillatory or rotational motion without the need to solve the equations of motion. This makes them a valuable tool in practical calculations. In the context of the EMRI modelling, we expect that the Hamiltonian frequencies will effectively characterize the periodicity of the orbit, in a manner analogous to that observed in EMRI systems with a non-spinning secondary, as described in Refs. [46, 53].
.2.1 Canonical formulation
@@
$\left(\tilde{t},r,\nu,\tilde{\ell},\varpi;\pi_{t},\tilde{\pi}_{r},\pi_{\nu},\pi_{\ell},\pi_{\varpi}\right)\,.$ (2)
@@
Table 1: Dictionary of special orbital and spin configurations. The alignment of spin is defined between the 4-spin $S^{a}_{\text{\tiny{TD}}}$ and the covariant angular momentum 4-vector $\mathcal{L}^{a}$, cf. Sec. I.3.4. It should be noted that planar motion is possible if and only if $S^{a}_{\text{\tiny{TD}}}$ is aligned with $\mathcal{L}^{a}$.
.2.2 Action-angle variables and Hamiltonian frequencies
The integrability of the Hamiltonian system in Eq. (3) directly implies the existence of action-angle variables for $\mathcal{P}$  [^1] We note that action-angle variables are denoted here with the angles $\vartheta^{i}$ first and the action $\mathcal{J}_{i}$ second. This is conventional in physics. We also use interchangeably “variable” and “coordinates”. $(\vartheta^{i},\mathcal{J}_{i})_{i\in\{1\ldots 5\}}\in\mathbb{T}^{5}\times\mathbb{R}^{5}$. Using the classical method based on a type-$2$ generating function [cf. Refs. [53, 46, 62]], the action coordinates (for bound orbits) can be constructed as follows  [^2] Action-angle coordinates are typically presented in the form of distinct patches, each covering a specific region of the phase space. In this study, we will focus on bound orbits.
@@
As a final note, we mention that there are several independent efforts to derive the action-angle variables for the spinning secondaries orbiting primary black holes; see, for example, Refs. [64, 65, 36, 66]. These works generalize our own in the sense that they are for primary Kerr black holes, not just for Schwarzschild primaries. They are also restricted versions of our own, in the sense that they either are concerned with unbound orbits, assume the weak-field, post-Newtonian gravity, or are based on the Marck tetrad formalism which is limited to the linear-in-spin approximation, while our framework [2, 1] can be extended to the quadratic-in-spin case.

### v3

@@ Abstract
This work constitutes the second part of a series of studies that aim to utilise tools from Hamiltonian mechanics to investigate the motion of an extended body in general relativity. The first part of this work [Refs. [1, 2]] constructed a ten-dimensional, covariant Hamiltonian framework encompassing all the linear-in-spin corrections from the secondary spin to the geodesic motion in arbitrary spacetime. This framework was proven to be integrable in the Schwarzschild and Kerr spacetimes, specifically. The present work translates this abstract integrability result into tangible applications for linear-in-spin Hamiltonian dynamics of a compact object in a Schwarzschild spacetime. In particular, a canonical system of coordinates is constructed explicitly, which exploits the spherical symmetry of the Schwarzschild spacetime. These coordinates are based on a relativistic generalization of the classical Andoyer variables of Newtonian rigid body motion. This canonical setup allows us to derive ready-to-use formulae for action-angle coordinates and gauge-invariant Hamiltonian frequencies, which automatically include all linear-in-spin effects. No external parameters or ad hoc choices are necessary, and the framework can be used to find complete solutions by quadrature of generic (bound or unbound), linear-in-spin orbits, including orbital inclination, precession and eccentricity, as well as spin precession. The efficacy of the formalism is demonstrated here in the context of circular orbits with arbitrary spin and orbital precession, with the results validated against known results in the literature.
@@
The main motivation of our efforts comes from extreme mass-ratio inspirals (EMRIs) – the gradual inspiral of a stellar-mass compact object (typically a neutron star or black hole or even an “exotic” compact object) into a supermassive black hole [3]. These binary inspirals are promising sources of gravitational waves for an interferometric detector in space (e.g., the Laser Interferometer Space Antenna [4, 5], as well as DECIGO [6], TianQin, and Taiji [7, 8, 9, 10]) and they will eventually provide an exquisite probe of the strongly gravitating region close to a black hole; for example, the source’s astrophysical parameters, such as the masses and the spin of the larger central black hole, can be measured with unparalleled precision [11, 12]. However, the gravitational-wave signals from EMRIs will be so intricate and sufficiently weak that they can only be extracted from the noisy detector data with modelled waveforms, that match the phase of the signal over its full duration (scaling like the inverse of the mass ratio) across all possible orbital and spin configurations around a central black hole [13]. As a result, the scientific promise described above will be attained only with a highly accurate and efficient model of relativistic two-body systems, built on the detailed understanding of their orbital dynamics in the (multidimensional) EMRI parameter space.
@@
In what follows, we give some more details about the relativistic two-body problem in the extreme mass-ratio regime, and motivate our work by emphasising the need for a Hamiltonian treatment of the secondary spin.
@@
Over the years, exact numerical relativity simulations [14, 15, 16, 17, 18, 19, 20], analytical approximation schemes [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31] and effective models [32, 33, 34, 35, 36, 37, 38, 39] have been devised to tackle the relativistic two-body problem [40], each specializing to some type of source, but also overlapping with the other in their respective domain of validity [41, 42, 43, 44, 45, 46]. A point mass in general relativity travels on a pre-defined geodesic in the curved geometry of spacetime. However, this feature is only an approximation in the test-particle limit, as revealed by multipolar expansion schemes to a small, extended, compact object (surrounded by a vacuum region)  [47, 48, 49, 50]. This modelling approach represents the extended object as a point particle but endowed with a finite number of multipoles, encompassing mass, linear and angular momentum, deformability coefficients, among others. The granularity of this description increases with the number of multipoles included, providing a more realistic description of the object.
@@
In this multipolar expansion approach to the two-body dynamics, the associated small parameter is the typical length scale of the object over the typical length scale of the background spacetime, which is therefore tailor-made for modelling asymmetric binaries, in particular EMRIs. At the “zeroth” order, the motion of a smaller compact object is well approximated by a test body moving along a geodesics of the fixed background spacetime. However, this is only a crude approximation. Additional effects arise and lead to corrections of this geodesic motion [49]. These include (i) conservative and dissipative self-field effects of the small object, owing to the non-linearity of general relativity (GR) and the emission of gravitational waves, and (ii) extended-body effects, due to the compact object not being exactly a point, but rather an extended object which can spin and deform. All these effects are important for measuring EMRIs by space-based gravitational-wave detectors, as all of them have an impact of quadratic order in the mass ratio in the waveform phase [25, 51], which is considered to be the minimal accuracy required for parameter estimation [13, 52].
@@
In the present paper (and its foundation [1, 2]), we leave the self-force effects aside (i.e. we neglect the back-reaction of the small object’s gravitational field), and solely focus on the leading order, linear-in-spin corrections from the secondary’s extended body to the geodesic motion. At this linear-in-spin order, although it is specific to GR, the equations that describe the evolution of linear and angular momenta along the particle’s worldline are purely kinematical (no body-specific multipolar forces or torques) and universal (all bodies move the same way) [50, 53]. Body-specific forces and torques only arise at the next quadrupolar order, with quadratic-in-spin effects [54].
@@
Analytically solving the relevant linear-in-spin evolution equations, now known as the Mathisson-Papapetrou-Tulczyjew-Dixon (MPTD) equations, on an astrophysically relevant Kerr black-hole background has been the subject of many past studies using a whole spectrum of methods. Among them, state-of-the-art methods use the powerful formalism of the Marck tetrad [55, 56] for the spin to solve the MPTD equations directly in a “perturbed geodesic” philosophy. This method is very natural given the quasi-geodesic nature of linear-in-spin dynamics, especially leading to purely analytical results for generic orbits in the Schwarzschild spacetime [57] and simple orbital configurations in Kerr [58, 59, 60, 51, 61]. A hybrid numerical/analytical result exists for more generic orbits [62, 63, 64, 65]. As such, there is a sense in which this Marck formalism offers a simpler implementation than other means of obtaining the solutions to the linear-in-spin equations of motion. Having the solution trajectory, however, is not the only way to understand the linear-in-spin dynamics of the MPTD system. Many other useful characterizations such as integrability of the system, in which the equations of motion can be solved by quadrature (as in, e.g., Refs. [66, 57]) for any orbital configurations, may require another approach to the MPTD problem.
@@
A Hamiltonian formulation just offers such an approach. In the test-mass approximation (i.e. with no spin), the powerful machinery of Hamiltonian mechanics has been already proven its value in the context of motion in curved spacetime. A non-exhausted list of examples includes (but not limited to) the Hamilton-Jacobi equation [67], action-angle variables [68, 69], the phase space analysis [70], Hamiltonian frequencies and resonances [68, 71], the first law of mechanics [72, 73, 74], the perturbation (KAM-type) theory [75, 76], Birkhoff normal forms (through Lie series) [77] symplectic integrators [78, 79]. From the mathematical point of view, these tools are indeed brought by the unique integrability of Kerr geodesic dynamics as a Hamiltonian system [80].
Nonetheless, we are nowhere near having a Hamiltonian understanding of these Kerr, linear-in-spin dynamics as thorough as that of the Kerr geodesics dynamics. The main issue with including the spin is that the treatment becomes necessarily perturbative, since the MPTD equations themselves are always truncated to some multipolar order, for physical consistency (see details in Sec. I.1.) This perturbative treatment makes ordinary Hamiltonian concepts such as “an integrable system” or the definition of action-angle variables perturbative as well. We shall follow this approach here as well, and refer to Ref. [81] for details on the mathematical well-posedness of these concepts. On top of this difficulty, the requirement for a spin-supplementary condition (SSC) when the spin is involved in the dynamics implies the existence of algebraic constraints in addition to the differential equations of motion. From a symplectic mechanics point of view, this means that one must reduce the original, non-constrained phase space, to a lower-dimensional one, where the SSC holds. Hamiltonian schemes including linear-in-spin effects exist, for example those constructed in Refs. [82, 83, 84, 85, 86] (see also Refs. [1] for detailed references). In particular, based on Ref. [86], the framework presented in Ref. [66] allowed a well-posed definition of frequencies of motion and (analytical) solutions for spherically symmetric backgrounds in Ref. [57, 87] as well as those for the Kerr background in Ref. [88] under the Tulzyjew-Dixon (TD) spin supplementary condition. It is, however, based on the Marck tetrad formalism in Ref. [55], which intrinsically relies on describing spinning trajectories as perturbation of (non-spinning) geodesics. In a similar vein, Refs. [85, 89] (and references therein) established the linear-in-spin integrability of the system under another spin supplementary condition, the Newton-Wigner one.
It is not trivial to extend these linear-in-spin formalisms to the quadratic-in-spin case: the Marck tetrad formulation relies on the parallel transport equation for the spin, which is not valid beyond linear order, while the Newton-Wigner spin supplementary condition does not lead to the existence of extra constants of motion, which are key to exploit the quasi-integrability of the dynamics. The present paper, and its foundation in Refs. [2, 1], aim at providing a rigorous treatment of the linear in spin dynamics, backed up by mathematical theorems on Poisson geometry, for the Hamiltonian analysis at linear order in spin, in a Schwarzschild background. It will provide a basis on which to extend the analysis to quadrupolar, quadratic-in-spin dynamics in subsequent works.
.1.1 This work: the canonical formalism in the Schwarzschild spacetime
@@
The classical multipolar expansion schemes (reviewed in Sec. I.1) assert that this object can be modelled as a point particle endowed with a linear momentum 1-form $p_{a}$ and spin tensor $S^{ab}$. Their evolution along the particle’s worldline $\mathscr{L}$, at linear order in $S^{ab}$ and under the Tulczyjew-Dixon (TD) spin supplementary condition (which fixes a unique representative worldline for the extended body), is given by the following set of (linearized) MPTD equations
@@
As previously outlined by one of us (PR) in Ref. [1] (hereafter referred to as “Paper I”) the algebraic-differential system in Eq. (1) can be cast into a covariant Hamiltonian formulation on a $10$ dimensional ($10$D) phase space $\mathcal{P}$. Most importantly, Paper I demonstrates that, when applied to a background Kerr spacetime, the resulting Hamiltonian system is integrable in the usual, Arnold-Liouville sense [cf Ref. [90, 91]].
@@
In the present paper we take on that endeavour, and construct canonical coordinates on $\mathcal{P}$ explicitly (see also Ref. [92] for other attempts in this direction). Based on these coordinates, we also derive a manifestly integrable Hamiltonian system which is effectively 1D system. This allows us to construct action-angle coordinates in a straightforward way. These action-angle coordinates permit the calculation of frequencies associated with oscillatory or rotational motion without the need to solve the equations of motion. This makes them a valuable tool in practical calculations.
It is important to acknowledge that there are several independent research projects aiming to derive action-angle variables for spinning secondaries orbiting a primary black hole; see, for example, Refs. [93, 66, 57, 94, 95]. These works generalize our own in the sense that they are for primary Kerr black holes, not just for Schwarzschild primaries. They are also restricted versions of our own, in the sense that they either are concerned with unbound orbits, assume the weak-field, post-Newtonian gravity, or are based on the Marck tetrad formalism which is limited to the linear-in-spin approximation, while our framework [2, 1] can be extended to the quadratic-in-spin case. In the context of EMRI modelling, we expect that the Hamiltonian frequencies will effectively characterize the periodicity of the orbit, in a manner analogous to that observed in EMRI systems with a non-spinning secondary, as described in Refs. [68, 75].
.2 Outline of the work
We shall provide in this section an overview of our paper and its main results. We begin in Sec. .2.1 by outlining the construction of the canonical coordinate of the physical phase space $\mathcal{P}$, followed by an examination of the associated canonical formalism of the Hamiltonian system on phase space $\mathcal{P}$ in Sec. .2.2. The action-angle coordinates and Hamilton frequencies of the system are presented in Sec. .2.3.
As previously stated, the present work is based on our general, covariant Hamiltonian framework for the dynamics of an extended (‘test’) body in curved spacetime. See already Sec. I a review of the framework presented in Paper I [1]; for a more detailed overview, readers may also consult the letter version of Paper I: Ref. [2].
.2.1 Brief overview of the construction
This will be the content of Sec. IIIII of the body of the paper. We outline the main points here.
Like previous studies on the covariant Hamiltonian formulation for the dynamics of spinning secondary bodies (e.g. Refs. [86, 96, 66]), our framework begins by casting the MPTD system in Eq. (1) into the most general $14$D Poisson manifold $\mathcal{M}$, which is degenerate. This Poisson system is mapped to a non-degenerate $12$D symplectic manifold $\mathcal{N}$, which does not respect the spin supplemental condition. This symplectic manifold is then projected onto a non-degenerate $10$D physical phase space $\mathcal{P}$, where the spin supplemental conditions are identically satisfied. For an illustration of this reduction process, see also FIG. 1, along with the accompanying discussion in Sec. I.2.
In parallel, the construction of the canonical coordinate  [^1] We shall use interchangeably “variable” and “coordinates” when there is no danger of confusion. is based on the generic coordinate $(x^{\alpha},p_{\beta},S^{\gamma\delta})$ associated with the $14$D Poisson manifold $\mathcal{M}$ (for details, see Sec. I.2). From this fundamental coordinate, the canonical coordinate $y_{\mathcal{N}}$ of the $12$D symplectic manifold $\mathcal{N}$ [cf. Eq. (20)] is derived as described in Paper I. It should be noted that these coordinates can be utilized in any curved background spacetime. However, due to this versatility, they may not always be the most optimal coordinates for a specific spacetime, especially if it has symmetry. This is the focus of our investigation in this paper.
Figure 1: The relevant phase space and its sub-manifolds for the covariant Hamiltonian formulation of the linear-in-spin system presented in Refs. [1, 2]. We use different sub-manifolds to lift the degeneracies associated with the existence of Casimir invariants $(\mathcal{M}\rightarrow\mathcal{N})$ and correctly implement the TD spin supplemental conditions $(\mathcal{N}\rightarrow\mathcal{T}\rightarrow\mathcal{P})$. Geodesics dynamics are recovered in the non-spinning limit, spanning the $8$D phase space $\mathcal{G}$.
In order to fully exploit the spherical symmetry inherent to Schwarzschild geometry, we put forth the canonical transformation of $y_{\mathcal{N}}\mapsto y_{\mathcal{N}}^{A}$ in Sec. II, which generalises the conventional orbital elements for Schwarzschild geodesics dynamics to those applicable to the dynamics of a spinning secondary. We call the resulting $12$ coordinates $y_{\mathcal{N}}^{A}$: the (relativistic) Andoyer variables, and they are presented in Eq. (66) of Sec. II.2,
The following Sec. III is dedicated to apply the theory of constrained Hamiltonian systems which enables the transition from $12$D symplectic manifold $\mathcal{N}$ to its $10$D physical submanifold $\mathcal{P}$, where the spin supplementary condition is held to be true. Because the Andoyer variables $y_{\mathcal{N}}^{A}$ inherited from $\mathcal{N}$ are no longer canonical on $\mathcal{P}$, not merely due to the differing dimensions of the phase spaces involved, the construction of canonical coordinate on $\mathcal{P}$ will proceed in two step. In Sec. III.1, we introduce the $10$ ‘complete’ (yet non-canonical) variables $y_{\mathcal{P}}^{C}$ [cf. Eq. (80)] as intermediate coordinates. In the second step in Sec. III.2, we then transform $y_{\mathcal{P}}^{C}$ to the final $10$ canonical coordinate: $y_{\mathcal{P}}^{F}$. This use of $y_{\mathcal{P}}^{C}$ as intermediate non-canonical variables is for purely pragmatic, in order to facilitate our search for the final form of $y_{\mathcal{P}}^{F}$.
In summary, we have the following reduction procedure, also illustrated in Eq. (2):
@@
It should be highlighted that all of the coordinate systems in Eq. (2) that emerged during the reduction process are connected to one another through one-to-one mappings. No additional assumptions have been made, apart from the linear-in-spin approximation. Therefore, it is a simple process to move between each of the variables.
Consequently, any given Hamiltonian associated with distinct variables ($y_{\mathcal{N}},y_{\mathcal{N}}^{\textrm{A}},y_{\mathcal{P}}^{\textrm{F}}$, etc.) and a different phase space ($\mathcal{N},\mathcal{P}$, etc.) can be derived by simply inserting the mapping in Eq. (2) into the original Hamiltonian $H_{\mathcal{M}}:\mathcal{M}\rightarrow\mathbb{R}$ of the initial $14$D Poisson system with $\mathcal{M}$, expressed in terms of the original variables $(x^{\alpha},p_{\beta},S^{\gamma\delta})$. To illustrate, for example, the final Hamiltonian $H_{c}:\mathcal{P}\rightarrow\mathbb{R}$ of the $10$D physical phase space $\mathcal{P}$ can be obtained schematically by $H_{c}:=H_{\mathcal{M}}[x^{\alpha}(y_{\mathcal{P}}^{\textrm{F}}),p_{\beta}(y_{\mathcal{P}}^{\textrm{F}}),S^{\gamma\delta}(y_{\mathcal{P}}^{\textrm{F}})]$, whose explicit expression will be shortly presented in Eq. (4) in Sec. .2.2.
The reader is directed to Table. 2 at the end of this introduction [cf. Sec. .4], which sets out the precise location of each mapping in explicit forms.
.2.2 Canonical formalism
The content of the present subsection, along with the following one, will be the content of Sec. IV of the main body of the paper. Again, we outline the main points here.
[…]
$y_{\mathcal{P}}^{F}:=\left(\tilde{t},r,\nu,\tilde{\ell},\varpi;\pi_{t},\tilde{\pi}_{r},\pi_{\nu},\pi_{\ell},\pi_{\varpi}\right)\,.$ (3)
@@
Table 1: Dictionary of special orbital and spin configurations. The alignment of spin is defined between the 4-spin $S^{a}_{\text{\tiny{TD}}}$ and the covariant angular momentum 4-vector $\mathcal{L}^{a}$ [cf. Sec. I.3.4]. It should be noted that planar motion is possible if and only if $S^{a}_{\text{\tiny{TD}}}$ is aligned with $\mathcal{L}^{a}$.
.2.3 Action-angle variables and Hamiltonian frequencies
The integrability of the Hamiltonian system in Eq. (4) directly implies the existence of generalised action-angle variables $(\vartheta^{i},\mathcal{J}_{i})_{i\in\{1\ldots 5\}}\in\mathbb{T}^{5}\times\mathbb{R}^{5}$  [^2] We note that action-angle variables are denoted here with the angles $\vartheta^{i}$ first and the action $\mathcal{J}_{i}$ second. This is conventional in physics., in the sense of Ref.[68, 75] to handle the usual, non-compact time direction. Using the classical method based on a type-$2$ generating function [cf. Refs. [75, 68, 91]], the action coordinates (for bound orbits) can be constructed as follows  [^3] Action-angle coordinates are typically presented in the form of distinct patches, each covering a specific region of the phase space. In this study, we will focus on bound orbits.